OpenStudy (anonymous):

find the difference of -8-(-11) would that answer me 3, -3, 19 or -19

OpenStudy (anonymous):

3

OpenStudy (anonymous):

-8 - - 11 = -8 + 11 = 3

OpenStudy (anonymous):

-8 minus a negative 11 would give you 3

OpenStudy (anonymous):

Find the sum -7 + (-2) would it be 5,-5,9 or -9

OpenStudy (anonymous):

-7+(-2)=-7-2=-9

OpenStudy (anonymous):

'--' = + '-+' or +-' = -